Find the integral. 786 ON 10083* * 08 0604x30 76 06 Oxato O D. 1 (0+x?) % +c Click to select your answer.
Answer #1

Given that I = 1-7X (9+x?js Ar=!(9+x?j*** Put 9+x? = u 7xⓇdx = du a = + = = (: u = 9+x) -5+1 +C=-- 49+x?)* 4 Therefore the co